OpenStudy (anonymous):

21 mariposa007 0Find the hypotenuse of a right triangle with legs of 9 inches and 12 inches= 12?

OpenStudy (anonymous):

\[a^2+b^2=h^2\] if they are sides of a right triangle therefore \[9^2+12^2=h^2\]

OpenStudy (anonymous):

81+144=195

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

seems unlikely

OpenStudy (anonymous):

\(81+144=?\)

OpenStudy (anonymous):

LOL sorry 225

OpenStudy (anonymous):

15 answer

OpenStudy (anonymous):

whew then since \(h^2=225\) you know \(h=\sqrt{225}=15\)

OpenStudy (anonymous):

which actually you knew at the start, since \(9,12,15\) is really just \(3,4,5\) in disguise